# MARS: Multi-Agent Robotic System with Multimodal Large Language Models for Assistive Intelligence

### v1

@@ Abstract
Multimodal large language models (MLLMs) have shown remarkable capabilities in cross-modal understanding and reasoning, offering new opportunities for intelligent assistive systems, yet existing systems still struggle with risk-aware planning, user personalization, and grounding language plans into executable skills in cluttered homes. We introduce MARS—a Multi-Agent Robotic System powered by MLLMs for assistive intelligence and designed for smart home robots supporting people with disabilities. The system integrates four agents: a visual perception agent for extracting semantic and spatial features from environment images, a risk assessment agent for identifying and prioritizing hazards, a planning agent for generating executable action sequences, and an evaluation agent for iterative optimization. By combining multimodal perception with hierarchical multi-agent decision-making, the framework enables adaptive, risk-aware, and personalized assistance in dynamic indoor environments. Experiments on multiple datasets demonstrate the superior overall performance of the proposed system in risk-aware planning and coordinated multi-agent execution compared with state-of-the-art multimodal models. The proposed approach also highlights the potential of collaborative AI for practical assistive scenarios and provides a generalizable methodology for deploying MLLM-enabled multi-agent systems in real-world environments.
@@
[…]
[inst2]organization=Chongqing University, Chongqing 401331, China
@@ 1 Introduction
To address these challenges, we propose a MLLM-based multi-agent collaborative system. For the assistive robotics domain, the input branch first preprocesses static images of the user’s surrounding environment and, by integrating depth and semantic features, extracts key objects/targets such as indoor layouts, object positions, and human presence. Based on this, a visual agent is constructed using an MLLM combined with models like CLIP, capable of recognizing and extracting environmental information such as obstacle types, locations, and passage widths, thereby satisfying perception requirements. Leveraging the perception output, the risk reasoning branch employs multi-agent collaboration mechanisms to analyze environmental information, assess potential hazards, and generate prioritized handling sequences. The planning and iterative branch then generates preliminary action sequences based on risk priority, which are evaluated by the evaluation agent to detect potential conflicts and risks, with feedback guiding subsequent adjustments. The system ultimately outputs robot-executable physical actions and path plans.
Overall, the contributions of this work are as follows:
@@ 3.1 Overall System Architecture
The architecture follows a closed-loop cycle of perception, reasoning, planning, evaluation, iteration, ensuring coordination and continuous feedback among agents. As shown in Figure 1, the system consists of an input layer, four core agent modules, and an output interface.
The input layer receives RGB images of the user’s surroundings, with optional depth maps or semantic segmentation results to enrich spatial and semantic understanding. The visual perception agent detects obstacles, furniture, users, and assistive devices, and evaluates pathway accessibility. Based on these outputs, the reasoning agent identifies risks and assigns severity and priority levels. The planning agent then generates executable action plans for robots, including obstacle removal, object adjustment, and path guidance. The feedback agent evaluates plan feasibility and safety, returning to the planning stage for revision when needed. Finally, the output interface delivers execution commands to the robot and provides alerts or guidance to the user or caregiver.
This closed-loop design integrates perception, reasoning, planning, and feedback into a unified process, enhancing system robustness in dynamic and complex environments.
@@ 3.2 Workflow
The system workflow consists of four stages: perception, risk analysis, planning, and iterative optimization, as shown in Algorithm 1. In the perception stage, the visual agent processes environmental images with multimodal models and detection algorithms to produce structured scene representations containing object categories, coordinates, and spatial relations. In the risk analysis stage, the reasoning agent detects hazards—such as blocked pathways, oversized furniture, or narrow passages—and prioritizes them by urgency and severity. The planning stage then generates preliminary action sequences or finite-state machines executable by robots, covering tasks such as obstacle clearance, furniture repositioning, and safe navigation.
During iterative optimization, the feedback agent verifies constraints on the plan, ensuring safety and feasibility. If conflicts or risks are detected, the system returns to the planning stage until a valid solution is produced. The final output includes both robot-executable commands and user-oriented feedback, ensuring operational safety while maintaining transparency in human–robot interaction.
@@ 3.3 Multi-Agent Collaboration Module Design
3.3.1 I. Visual Perception Agent
@@
3.3.2 II. Risk Assessment and Reasoning Agent
@@
In the risk identification stage, the agent matches structured elements in $\mathcal{D}_{\text{Agent1}}$ against predefined risk categories and localizes their positions. Referring to the typical categories in (Table LABEL:tab:risk), each scene element is systematically checked. For passage-related risks, the agent analyzes passage width and obstacle distribution against safety thresholds. Let the minimum wheelchair-accessible passage width be $\tau_{w}(e.g.,\tau_{w}=0.8\,\text{m})$. If the actual passage width satisfies $width_{c}<\tau_{w}$, the case is flagged as a “narrow passage risk.” If the obstacle set in the passage is nonempty, i.e., $obstacle_{c}\neq\emptyset$, the obstacle ratio $r_{obs}$ is computed as
@@
3.3.4 IV. Evaluation and Optimization Agent
The evaluation and optimization agent is responsible for assessing the quality of planning schemes and guiding their refinement, thereby improving overall performance through multidimensional evaluation and a closed-loop feedback mechanism. Unlike traditional single-mode feasibility validation, this agent adopts a predefined four-dimensional evaluation framework that enables a holistic assessment of initial plans, balancing user needs with system robustness. Its input is the preliminary action sequence $\mathcal{A}=[a_{1},a_{2},...,a_{k}]$ generated by the planning and task allocation agent. The evaluation is conducted along four dimensions defined in Table 2, each associated with specific criteria (e.g., Assistance UX covers clarity of interaction, usability, and trust; Task Efficiency addresses time efficiency, resource utilization, and robustness), forming a multi-perspective assessment network.
@@ 5 Discussion and Conclusion
Peiyan Zhong: Conducted ablation analyses; contributed to the writing and revisions; prepared the figures.

### v2

@@ Abstract
Multimodal large language models (MLLMs) have shown remarkable capabilities in cross-modal understanding and reasoning, offering new opportunities for intelligent assistive systems, yet existing systems still struggle with risk-aware planning, user personalization, and grounding language plans into executable skills in cluttered homes. We introduce MARS, a Multi-Agent Robotic System powered by MLLMs for assistive intelligence and designed for smart home robots supporting people with disabilities. The system integrates four agents: a visual perception agent for extracting semantic and spatial features from environment images, a risk assessment agent for identifying and prioritizing hazards, a planning agent for generating executable action sequences, and an evaluation agent for iterative optimization. By combining multimodal perception with hierarchical multi-agent decision-making, the framework enables adaptive, risk-aware, and personalized assistance in dynamic indoor environments. Experiments on multiple datasets demonstrate the superior overall performance of the proposed system in risk-aware planning and coordinated multi-agent execution compared with state-of-the-art multimodal models. The proposed approach also highlights the potential of collaborative AI for practical assistive scenarios and provides a generalizable methodology for deploying MLLM-enabled multi-agent systems in real-world environments.
@@ 1 Introduction
To address these challenges, we propose a MLLM-based multi-agent collaborative system. For the assistive robotics domain, the input branch first preprocesses static images of the user’s surrounding environment and, by integrating depth and semantic features, extracts key objects/targets such as indoor layouts, object positions, and human presence. Based on this, a visual agent is constructed using an MLLM combined with models like CLIP, capable of recognizing and extracting environmental information such as obstacle types, locations, and passage widths, thereby satisfying perception requirements. Leveraging the perception output, the risk reasoning branch employs multi-agent collaboration mechanisms to analyze environmental information, assess potential hazards, and generate prioritized handling sequences. The planning and iterative branch then generates preliminary action sequences based on risk priority, which are evaluated by the evaluation agent to detect potential conflicts and risks, with feedback guiding subsequent adjustments. The system ultimately outputs robot-executable physical actions and path plans. Overall, the contributions of this work are as follows:
@@ 3.1 Overall System Architecture
The architecture follows a closed-loop cycle of perception, reasoning, planning, evaluation, iteration, ensuring coordination and continuous feedback among agents. As shown in Figure 1, the system is architected as a sequential pipeline comprising an input transduction layer, four core cognitive agent modules operating in a coordinated manner, and an output actuation interface.
The input layer is tasked with ingesting RGB images of the ambient environment, which can be optionally augmented with depth maps or precomputed semantic segmentation masks to improve the understanding of the geometric and semantic scenes of the system. The visual perception agent detects obstacles, furniture, users, and assistive devices, and evaluates pathway accessibility. Based on these outputs, the reasoning agent identifies risks and assigns severity and priority levels. The planning agent then generates executable action plans for robots, including obstacle removal, object adjustment, and path guidance. The feedback agent evaluates plan feasibility and safety, returning to the planning stage for revision when needed. Finally, the output interface delivers execution commands to the robot and provides alerts or guidance to the user or caregiver. This closed-loop design integrates perception, reasoning, planning, and feedback into a unified process, enhancing system robustness in dynamic and complex environments.
@@ 3.2 Workflow
The system workflow consists of four stages: perception, risk analysis, planning, and iterative optimization, as shown in Algorithm 1. In the perception stage, the visual agent processes environmental images with multimodal models and detection algorithms to produce structured scene representations containing object categories, coordinates, and spatial relations. In the risk analysis stage, the reasoning agent detects hazards, such as blocked pathways, oversized furniture, or narrow passages, and prioritizes them by urgency and severity. The planning stage then generates preliminary action sequences or finite state machines executable by robots, covering tasks such as obstacle clearing, furniture repositioning, and safe navigation.
The planning evaluation loop constitutes a generate and test optimization algorithm. The Planning Agent acts as a proposal generator, while the Evaluation Agent implements a multi-dimensional scoring function (defined by $S_{1},S_{2},S_{3},S_{4}$}). The loop’s termination condition ($\bar{S}$ ¿ threshold) and the heuristic feedback rule (suggestions based on $d_{min}$) together form a lightweight, heuristic search procedure that iteratively refines the action sequence $\mathcal{A}$ towards $\mathcal{A}^{*}$, ensuring Pareto like improvements across competing objectives. The final output includes both robot executable commands and user oriented feedback, ensuring operational safety while maintaining transparency in human–robot interaction.
@@ 3.3 Multi-Agent Collaboration Module Design
3.3.1 Visual Perception Agent
@@
3.3.2 Risk Assessment and Reasoning Agent
@@
The risk identification phase operates as a pattern matching and instantiation process. The agent performs a subgraph isomorphism check between the structured scene graph implicit in $\mathcal{D}{\text{Agent1}}$ and a library of predefined risk schema templates. Successful matches instantiate specific risk instances, which are then spatially localized within the coordinate frame of $I{\text{RGB}}$. Referring to the typical categories in (Table LABEL:tab:risk), each scene element is systematically checked. For passage-related risks, the agent analyzes passage width and obstacle distribution against safety thresholds. Let the minimum wheelchair-accessible passage width be $\tau_{w}(e.g.,\tau_{w}=0.8\,\text{m})$. If the actual passage width satisfies $width_{c}<\tau_{w}$, the case is flagged as a “narrow passage risk.” If the obstacle set in the passage is nonempty, i.e., $obstacle_{c}\neq\emptyset$, the obstacle ratio $r_{obs}$ is computed as
@@ 3.3.3 Planning Agent
3.3.4 Evaluation and Optimization Agent
The evaluation and optimization agent is responsible for assessing the quality of planning schemes and guiding their refinement, thereby improving overall performance through multidimensional evaluation and a closed-loop feedback mechanism. Departing from conventional unimodal feasibility verification, this agent implements a multi-criteria decision analysis (MCDA) framework predicated on four orthogonal evaluation dimensions. This framework facilitates a Pareto-optimal trade-off analysis among competing objectives, such as user-centric experience metrics and system-level robustness guarantees, during the holistic assessment of candidate plans. Its input is the preliminary action sequence $\mathcal{A}=[a_{1},a_{2},...,a_{k}]$ generated by the planning and task allocation agent. The evaluation is conducted along four dimensions defined in Table 2, each associated with specific criteria (e.g., Assistance UX covers clarity of interaction, usability, and trust; Task Efficiency addresses time efficiency, resource utilization, and robustness), forming a multi-perspective assessment network.